OpenStudy (anonymous):

4e-8= 32 how do i get 4e alone?

OpenStudy (anonymous):

you add 8 to both sides. 4e-8=32 +8 +8 4e=40

OpenStudy (anonymous):

did you mean to solve for e? or to just get 4e alone on one side?

OpenStudy (anonymous):

so i put 4e -8 + 8 =32 + 8 ?

OpenStudy (anonymous):

?